OpenStudy (anonymous):

Which two equations are equivalent? A. y = (x + 4)2 y = x2 + 16 B. y = (x – 3)2 y = x2 – 6x + 9 C. y = (x + 2)2 y = x2 + 2x + 2 D. y = (x – 6)2 y = x2 – 12x – 36

OpenStudy (anonymous):

Expand everything.

OpenStudy (anonymous):

i assume that is squared?

OpenStudy (anonymous):

if so the FOIL

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

Are you sure? Double check and multiply carefully.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

I think that is the equation A. .. do not know whether it is right ...

OpenStudy (anonymous):

ugh i have no clue how to do this

OpenStudy (anonymous):

A. (x+4)(x+4) FOIL.......... x^2+4x+4x+x^2=x^4+8x

OpenStudy (haseeb96):

B is the correct answer and i am fully sure

OpenStudy (anonymous):

oooo

OpenStudy (anonymous):

B. (x-3)(x-3) FOIL........... x^2-3x-3x-9=x^2-6x-9

OpenStudy (anonymous):

Thank You Everyone

OpenStudy (haseeb96):